OpenStudy (vshiroky):

Last week David Wright had a batting average of .350. After the next night's game, Wright had 3 hits in 5 at bats, his average increased to .352. Determine the number of at bats and his Wright has altogether.

OpenStudy (anonymous):

.350 is David's cumulative average.

OpenStudy (anonymous):

If you were to add the 3/5 result to compute the new average, the expression would be

OpenStudy (vshiroky):

I'm not sure what to do..

OpenStudy (anonymous):

Let x be the number of hits before the game, and y be the number of bats before the game.

OpenStudy (anonymous):

Then, you have two equations.

OpenStudy (anonymous):

\[\frac{ x }{ y } = .350\]

OpenStudy (anonymous):

\[\frac{ x+3 }{ y+5 } = .352\]

OpenStudy (anonymous):

Batting average = #hits/#bats

OpenStudy (anonymous):

Take the two equations above and solve for x, y.

OpenStudy (vshiroky):

how do i solve x&y?

OpenStudy (vshiroky):

@ganeshie8

OpenStudy (vshiroky):

Part 2: At different times in the season, when David Wright had 3 hits for 5 times at bat, his batting average jumped from .350 to .351, from .350 to .355. Explain why the increase is not the same amount each time.

OpenStudy (wolf1728):

vshiroky - Let's solve the first question okay? (x+3) / (y+5) = .352 we can multiply that x + 3 = .352y + 1.76 and x = .350 y

OpenStudy (wolf1728):

We can rearrange those 2 equations A) x -.352y = -1.24 B) x -350y =0 Can you solve that now?

OpenStudy (wolf1728):

Hint multiply equation B by -1

OpenStudy (vshiroky):

625

OpenStudy (wolf1728):

A) x -.352y = -1.24 Okay, I'LL multiply equation B by -1 B) -x +.350y = 0 Now, add both equations -.002y = -1.24 y = 620

OpenStudy (wolf1728):

x = .350y x = 217

OpenStudy (vshiroky):

Add 217 & 620?

OpenStudy (wolf1728):

SO, when the batting average was .350 it consisted of 217 hits for 620 at bats The batting average is computed as: hits / at bats 217 / 620 = .350

OpenStudy (vshiroky):

Do I add the 3 to 217? For the total bats to be 220

OpenStudy (vshiroky):

It wants to know at bat and hits all together I think

OpenStudy (wolf1728):

That would be TOTAL HITS = 220 Then add 5 to the at bats making that equal 625 220 / 625 = 0.352

OpenStudy (vshiroky):

So 845 would be my final answer right? To know at bat and hits together?

OpenStudy (wolf1728):

Now you know the hits and at bats for the .350 average AND the hits and at bats for the .352 average I think that answers the first question WAY up at the top of the page. NO your answer is not 845 please look at the answers again

OpenStudy (vshiroky):

I thought to get the # of times he bat and hit would be added but maybe 405 cause I'd have to minus the 2

OpenStudy (vshiroky):

Am I still way off @wolf1728

OpenStudy (vshiroky):

My phone doesn't load to fast. I had to switch off the laptop sorry

OpenStudy (wolf1728):

Basically, you have to understand how the batting avaerage works. if you have 4 hits for 10 at bats what is your batting avaerage?

OpenStudy (vshiroky):

.4

OpenStudy (vshiroky):

.363636

OpenStudy (wolf1728):

CORRECT or .400 as it is usually reported Now, let's say at your NEXT at bat you do NOT get a hit - can you figure out the batting average then? Basically what is the "hits" what is the "at bats" and what is the average?

OpenStudy (wolf1728):

okay the hits = 4 the at bats = 11 and the avaerage = 4/11 = 0.3636363636 OR .364 as it is usually reported

OpenStudy (wolf1728):

So now do you see how to report the hits and the at bats in the original problem?

OpenStudy (vshiroky):

Hits 220 At bats 625

OpenStudy (wolf1728):

Okay, yes Actually, in the first, part it originally asked for the hits and at bats for the .350 average 217 hits 620 at bats

OpenStudy (vshiroky):

Ok

OpenStudy (wolf1728):

Do you have a little bit better idea now?

OpenStudy (vshiroky):

Yes I think so. Thank you.

OpenStudy (wolf1728):

Okay - glad to have helped out :-)

OpenStudy (vshiroky):

I can't medal from my phone but I will from the computer tomorrow